A train leaves Toronto for Montreal at the same time as another train leaves Montreal for Toronto. The cities are 500km apart. The trains pass each other 2h later. The train from Montreal is traveling 50km/h faster than the one from Toronto. At what distance away from Toronto do the trains pass each other?
I do not understand as to how I should set up each equation. Solving is easy, but setting up the equations is hard.
What I know:
Toronto Montreal
D=500km D=500km
S=x S=x+50

Answer by lwsshak3(11628)

let x = distance from Montreal to the point where the trains pass each other after two hours.
then, 500-x =distance from Toronto to the point where the trains pass each other after two hours.
speed = distance ÷ time
speed of Montreal train = x/2
speed of Toronto train = (500-x)/2
The Montreal train is 50 km/hour faster
x/2 -(500-x)/2=50
x - 500+x=100
2x = 600
x=300
500-x = 200
ans: the distance away from Toronto as the trains pass each other is 200 km